Use the quadratic formula:
2x^2-5x=3
Can anyone possibly help me understand this and how to write the answer in the correct terms. Thank you so much !

Answer by checkley71(8403)

2x^2-5x-3=0.
THE QUADRATIC FORMULA IS
X=(-B+-SQRT[B^2-4AC])/2A FOR A FORMULA AX^2+BX+C=0
X=(5+-SQR[-5^2-4*2*-3])/2*2
X=(5+-SQRT[25+24])/4
X=(5+-SQRT1)/4
X=5+-1)/4
X=5+1/4
X=6/4
X=1.5 SOLUTION
X=(5-1)/4
X=4/4
X=1 SOLUTION
###################################################
HAPPY NEW YEAR